Any help on how to solve for x in this problem would be welcomed. Thank you for your time and knowledge.
Solve the equation for x.
ax^2 - (2a + 1)x + (a + 1) = 0 (a > 0)

Answer by mananth(16946)

ax^2 - (2a + 1)x + (a + 1) = 0 (a > 0)
ax^2-(2a+1)x +(a+1) = 0
ax^2 - 2ax - x + a +1 = 0
a(x^2 - 2x + 1) - (x - 1) = 0
a(x - 1)^2 - (x - 1) = 0
(x-1)(a(x-1) - 1) = 0
(x-1)(ax - (a+1)) = 0
a(x-1)(x - (a+1)/a) = 0
x=1 or x=(a+1)/a